# Faster Gradient Methods for Highly-smooth Stochastic Bilevel Optimization

### v1

@@ Appendix A Notations for Tensors
$\begin{cases}\mathbb{E}\left\|\Phi_{t}-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \mathbb{E}\left\|\Phi_{t}-\sum_{j=0}^{p}\beta_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
@@
$\begin{cases}\left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=0}^{p}\beta_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
@@
$\begin{split}\begin{cases}\|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=-p/2,\cdots,p/2,&\quad p\text{ is even};\\[5.69054pt] \|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=0,\cdots,p,&\quad p\text{ is odd}.\end{cases}\end{split}$ (17)
@@
We prove our lower bound for stochastic nonconvex-strongly-convex bilevel optimization via a reduction to the lower bound for stochastic single-level nonconvex optimization [3]. To state their lower bound, we first need to introduce the function class, oracle class, algorithm class, and the complexity measures.
@@
All the above definitions are merely restatements of [3, Section 2]. Although Definition E.4 uses the definition of distributional complexity, by Yao’s minimax principle is also a lower bound for the worst-case complexity. Now, we recall the construction in [3] for proving the $\Omega(\epsilon^{-4})$ lower bound. Formally, we define the randomized function
@@
where $\bar{L}_{1}=155$, $\beta=4\bar{L}_{1}\epsilon/L_{1}$, $\rho:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}^{T}$ is $\rho({\bm{x}})={\bm{x}}\big{/}\sqrt{1+\|{\bm{x}}\|^{2}/R^{2}}$, $R=230\sqrt{T}$, $\lambda=1/5$, and $f_{T}:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}$ is the nonconvex hard instance introduced by Carmon et al. [6]:
@@
For the hard instance in Eq. (19), Arjevani et al. [3] further defined the stochastic gradient estimator $F_{\bm{U}}$ as
@@
where ${\rm prog}_{\alpha}(x)=\max\{i\geq 0\mid|x_{i}|>\alpha\}$ and $\gamma=\min\{(46\epsilon)^{2}/\sigma^{2},1\}$. For the above construction, Arjevani et al. [3] showed the following lower bound.
Theorem E.1 ([3, Theorem 3]).
@@ Appendix F The F2SA-2 Algorithm
We present the realization of F2SA-$p$ when $p=2$ in Algorithm 3 to further compare its procedure with the original F2SA algorithm. Let $\lambda=1/\nu$. We can observe that F2SA [28, 10] solves the following asymmetric penalty problem
@@
while F2SA-2 solved the following symmetric penalty problem:
@@
3:  Sample random i.i.d indexes $(\xi^{y},\zeta^{y})$, $(\xi^{z},\zeta^{z})$, and $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
4:  ${\bm{y}}_{t}^{0}={\bm{y}}_{t},~{\bm{z}}_{t}^{0}={\bm{z}}_{t}$
5:  for $k=0,1,\cdots,K-1$
@@
10:  $\Phi_{t}=\dfrac{1}{2}\sum_{i=1}^{S}\left(F_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};{\xi_{i}^{x}})+F_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\xi_{i}^{x})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};\zeta_{i}^{x})-G_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\zeta_{i}^{x})}{\nu}\right)$
11:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
12:end for

### v2

@@ Appendix A Notations for Tensors
$\begin{cases}\mathbb{E}\left\|\Phi_{t}-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \mathbb{E}\left\|\Phi_{t}-\sum_{j=-(p-1)/2}^{(p+1)/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{t+1}^{j})}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
@@
$\begin{cases}\left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=-p/2}^{p/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is even};\\[14.22636pt] \left\|\nabla\varphi({\bm{x}}_{t})-\sum_{j=-(p-1)/2}^{(p+1)/2}\alpha_{j}\left(j\nabla_{x}f({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))+\dfrac{\nabla_{x}g({\bm{x}}_{t},{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t}))}{\nu}\right)\right\|={\mathcal{O}}(\epsilon),&\quad p\text{ is odd}.\end{cases}$
@@
$\begin{split}\begin{cases}\|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=-p/2,\cdots,p/2,&\quad p\text{ is even};\\[5.69054pt] \|{\bm{y}}_{t+1}^{j}-{\bm{y}}_{j\nu}^{*}({\bm{x}}_{t})\|={\mathcal{O}}(\nu\epsilon/L_{1}),~~\forall j=-(p-1)/2,\cdots,(p+1)/2,&\quad p\text{ is odd}.\end{cases}\end{split}$ (18)
@@
We prove our lower bound for stochastic nonconvex-strongly-convex bilevel optimization via a reduction to the lower bound for stochastic single-level nonconvex optimization (Arjevani et al., 2023). To state their lower bound, we first need to introduce the function class, oracle class, algorithm class, and the complexity measures.
@@
All the above definitions are merely restatements of (Arjevani et al., 2023, Section 2). Although Definition E.4 uses the definition of distributional complexity, by Yao’s minimax principle is also a lower bound for the worst-case complexity. Now, we recall the construction in (Arjevani et al., 2023) for proving the $\Omega(\epsilon^{-4})$ lower bound. Formally, we define the randomized function
@@
where $\bar{L}_{1}=155$, $\beta=4\bar{L}_{1}\epsilon/L_{1}$, $\rho:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}^{T}$ is $\rho({\bm{x}})={\bm{x}}\big/\sqrt{1+\|{\bm{x}}\|^{2}/R^{2}}$, $R=230\sqrt{T}$, $\lambda=1/5$, and $f_{T}:{\mathbb{R}}^{T}\rightarrow{\mathbb{R}}$ is the nonconvex hard instance introduced by Carmon et al. (2020):
@@
For the hard instance in Eq. (20), Arjevani et al. (2023) further defined the stochastic gradient estimator $F_{\bm{U}}$ as
@@
where ${\rm prog}_{\alpha}(x)=\max\{i\geq 0\mid|x_{i}|>\alpha\}$ and $\gamma=\min\{(46\epsilon)^{2}/\sigma^{2},1\}$. For the above construction, Arjevani et al. (2023) showed the following lower bound.
Theorem E.1 ((Arjevani et al., 2023, Theorem 3)).
@@ Appendix F The F2SA-2 Algorithm
We present the realization of F2SA-$p$ when $p=2$ in Algorithm 3 to further compare its procedure with the original F2SA algorithm. Let $\lambda=1/\nu$. We can observe that F2SA (Kwon et al., 2023; Chen et al., 2025b) solves the following asymmetric penalty problem
@@
As an improved method, the F2SA-2 solved the following symmetric penalty problem:
@@
3:  ${\bm{y}}_{t}^{0}={\bm{y}}_{t},~{\bm{z}}_{t}^{0}={\bm{z}}_{t}$
4:  for $k=0,1,\cdots,K-1$
5:   Sample random i.i.d indexes $(\xi^{y},\zeta^{y})$ and $(\xi^{z},\zeta^{z})$.
@@
10:  Sample random i.i.d indexes $\{({\xi_{i}^{x}},{\zeta_{i}^{x}})\}_{i=1}^{S}.$
11:  $\Phi_{t}=\dfrac{1}{2}\sum_{i=1}^{S}\left(F_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};{\xi_{i}^{x}})+F_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\xi_{i}^{x})+\dfrac{G_{x}({\bm{x}}_{t},{\bm{y}}_{t+1};\zeta_{i}^{x})-G_{x}({\bm{x}}_{t},{\bm{z}}_{t+1};\zeta_{i}^{x})}{\nu}\right)$
12:  ${\bm{x}}_{t+1}={\bm{x}}_{t}-\eta_{x}{\Phi_{t}}/{\|\Phi_{t}\|}$
13:end for
Appendix G Additional Experiments
This section provides additional experiments on finding the optimal per-parameter regularization of a 5-layer MLP with ReLU activation and the hidden layer size of $500$. Following the notation in Example 2.2, we let ${\bm{x}}\in{\mathbb{R}}^{d}$ parameterize the regularization matrix via ${\bm{W}}_{\bm{x}}={\rm diag}(\exp({\bm{x}}))$. We also let $\ell_{\rm val}$ and $\ell_{\rm tr}$ be the logistic loss of the network prediction on the validation set and training set, respectively. The problem to solve has the same formulation as Example 2.2, as restated below:
$\min_{{\bm{x}}\in{\mathbb{R}}^{d}}\ell_{\rm val}({\bm{y}}),\quad{\rm s.t.}\quad{\bm{y}}\in\arg\min_{{\bm{y}}\in{\mathbb{R}}^{d}}\ell_{\rm tr}({\bm{y}})+{\bm{y}}^{\top}{\bm{W}}_{{\bm{x}}}{\bm{y}}.$ (22)
The difference between Example 2.2 is that now the problem is nonsmooth nonconvex due to the use of the MLP model. We present the experiment results in Figure 2.
Figure 2: Performances of different algorithms on Problem (22) with an MLP model.
Appendix H Use of Large Language Models
Large language models were used to help calculate the coefficient $\alpha_{0}$ when $p$ is odd in Lemma 3.1, and to refine wording and correct grammatical errors in parts of the paper.